OpenStudy (stephanos100):

PLSS HELP!!! MEDALS!!! Find the exact value of cos(4pi/3): I know no that it is -1/2, but my original answer was -sqrt3/2, so I guess I just solved for the wrong angle. Can someone help me with this so I don't make this mistake again? Thanks!

OpenStudy (xguardians):

First, do you know how to convert radians to degrees?

OpenStudy (xguardians):

4*180/3 = 240 degrees cos(240)

OpenStudy (xguardians):

Then you have to reference that to the x-axis, which the closest being 180 degrees. Therefore, 240-180 = 60 cos(60)

OpenStudy (xguardians):

Then, hopefully, you have what I call it in school, the "chart of death" memorized, where cos(60) = -1/2

OpenStudy (xguardians):

Well, cos(60) = .5, but you have to remember, since it is in the third quadrant, it will be negative.

OpenStudy (kittiwitti1):

For problems like these, I would say you ask yourself these quesetions: What's the reference angle for \(\frac{4\pi}{3}\)? What is the x-value on the Unit Circle of that reference angle? (Hint: the reference angle is \(\frac{\pi}{3}\))

OpenStudy (stephanos100):

Oh okay right I forgot about the reference angle thingamajig. I was finding the angle from the terminal arm to the y axis but I should have been finding the angle from the terminal arm to the x axis. Thanks so much @kittiwitti1 !

OpenStudy (kittiwitti1):

No problem ! :) I use the Left-Hand Rule (LHR) to remember the sin and cos values for the reference angles, which are \(\frac{\pi}{6},\frac{\pi}{4}\text{ and }\frac{\pi}{3}\), or 30, 45 and 60 degrees, respectively.